OpenStudy (anonymous):

In a 500m race "A" beats "B" by 5seconds or 100 m.They decided to run another race and this time "A" gives "B" a head start of 200m.If "A" speed is twice his previous speed and B speed is one half of his previous speed,how far from the start should the winning post be so that they finish at same time?

OpenStudy (anonymous):

I dont know if this is right, i might be interpreting the problem wrong >.< anywhos, im thinking since A beat B by 5 secs, which was 100 m, then B's velocity is 100/5 = 20 m/s Then, we know he ran the 500m in 5 more secs than A did so: \[D = r_b(t_a+5) \iff 500 = 20(t_a+5) \iff t_a = 20 s\] which gives A a rate of: \[500 = r_a(20) \iff r_a = 25 m/s\] Still working out the rest.

OpenStudy (anonymous):

So now when they are running the second time: \[r_{a'} = 2r_a = 50 m/s, r_{b'} = \frac{1}{2}r_b = 10 m/s\] we have the distance is going to be: \[d = r_{a'}t\iff d = 50t, d = r_{b'}t+200 \iff d = 10t+200\] Solving this system gives: d = 250 m.

OpenStudy (anonymous):

Ans : 500 m

OpenStudy (anonymous):

wut? >.> lol let me look over junks again.

OpenStudy (anonymous):

i dont see how that answer makes sense =/ in the first scenario, A has to be running 25 m/s, and B has to be running 20 m/s in order for A to win by 5 secs (500m dash). So in the second race, A will be running at 50 m/s, and B will be running at 10 m/s if the answer was 500 m, A can run that in 10 secs, B would take 30 secs (he only has to run 300 m)

OpenStudy (anonymous):

In the answer 250m, it would take A 5 secs to run it, and B 5 secs (he only has to run 50 m)

OpenStudy (anonymous):

I just see the answer from answer key it is given in the book as 500m .I m not able to solve it.

OpenStudy (anonymous):

i must be interpreting the problem wrong then. It has to do something with how im calculating their speeds in the first race.